(4-2i)^2 = 12-16i how?
Answer by Alan3354(69443)

(4-2i)^2 = 12-16i how?
--------------
(4-2i)*(4-2i) = 4*(4-2i) - 2i*(4-2i)
= 16-8i - 8i+4i^2
= 16 - 16i + 4i^2
= 16 - 16i - 4
= 12 - 16i